I am trying to work this equation and can't seem to get the answer indicated in my textbook.
(4/5)x-(1/5)>=-5

Found 2 solutions by Alan3354, mananth:
Answer by Alan3354(69443)

(4/5)x-(1/5)>=-5
Multiply by 5
4x - 1 = -25
4x = -24
x = -6

Answer by mananth(16949)

(4/5)x-(1/5)>=-5
..
(4/5)x >=-5+1/5
(4/5)x>= -24/5
5*4x>=5*-24
4x>=-24
/4
x>=-6
....
m.ananth@hotmail.ca